# Unveiling Multiple Descents in Unsupervised Autoencoders

### v2

@@ Abstract
The phenomenon of double descent has challenged the traditional bias-variance trade-off in supervised learning but remains unexplored in unsupervised learning, with some studies arguing for its absence. In this study, we first demonstrate analytically that double descent does not occur in linear unsupervised autoencoders (AEs). In contrast, we show for the first time that both double and triple descent can be observed with nonlinear AEs across various data models and architectural designs. We examine the effects of partial sample and feature noise and highlight the importance of bottleneck size in influencing the double descent curve. Through extensive experiments on both synthetic and real datasets, we uncover model-wise, epoch-wise, and sample-wise double descent across several data types and architectures. Our findings indicate that over-parameterized models not only improve reconstruction but also enhance performance in downstream tasks such as anomaly detection and domain adaptation, highlighting their practical value in complex real-world scenarios.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Demonstration of double descent phenomenon with unsupervised AEs for the ”sample noise” scenario (Subsection 3.1). We present the test loss for varying epochs and hidden layer sizes.
In this study, we use AEs to explore double descent and its impact on key unsupervised tasks like domain adaptation, anomaly detection, and robustness to noisy data. AEs are widely used in unsupervised tasks such as denoising (Vincent et al., 2008, 2010), manifold learning (Wang et al., 2014; Duque et al., 2020), clustering (Song et al., 2013; Yang et al., 2019), anomaly detection (Sakurada & Yairi, 2014; Zhou & Paffenroth, 2017), feature selection (Han et al., 2018; Gong et al., 2022), domain adaptation (Deng et al., 2014; Yang et al., 2021), segmentation (Myronenko, 2019; Baur et al., 2021), and generative modeling (Kingma, 2013; Doersch, 2016), making them a prominent use case when studying double descent in unsupervised learning.
We start by analyzing linear AEs, whose lack of nonlinear activation functions allows for theoretical exploration. We prove and empirically confirm that double descent does not occur in linear unsupervised AEs. We then investigate nonlinear AEs and find that both double and triple descent emerge across different data models and architectures.
We provide extensive empirical evidence demonstrating the phenomenon in unsupervised nonlinear AEs trained on contaminated data, with ”memorization” causing overfitting and performance degradation. However, we find that over-parameterized models can extract the true signal, leading to a second descent and improved performance.
@@
These findings provide critical insights into the dynamics of unsupervised learning and the interplay between model capacity, noise, and performance. The phenomenon is illustrated in Figure 1, where we show the classic bias-variance tradeoff, the critical regime, and the second descent induced by the over-parametrization of unsupervised AEs.
Our findings have important implications for key tasks in unsupervised learning. Specifically, we show that over-parameterized models adapt better to target domains when trained on source domains despite distributional shifts. We also reveal non-monotonic performance in anomaly detection as model complexity increases, underscoring the practical relevance of our results, particularly when addressing outliers and domain shifts.
@@
(a) To the best of our knowledge, this is the first demonstration of model, epoch, and sample-wise double descent in a fully unsupervised setting. We analyze and contrast the behavior of linear and nonlinear AEs through analytical and empirical evaluations, highlighting their distinct dynamics.
@@ 2 Related Work
In this work, we study AEs’ unsupervised objective of minimizing reconstruction error, $\|\Phi({\boldsymbol{x}})-{\boldsymbol{x}}\|^{2}_{2}$, where ${\boldsymbol{x}}$ is the input and $\Phi({\boldsymbol{x}})$ is the AE’s output. This learning framework, which has not previously shown double descent, falls outside the settings explored in prior studies. We find that linear AEs lack double descent, while nonlinear AEs show double and triple descent, underscoring the role of nonlinearity. Our analysis spans model, epoch, and sample levels, exploring the impact of bottleneck size, hidden layers, noise magnitude, and sample contamination. We also demonstrate double descent in real-world unsupervised scenarios like domain shifts, anomalies, and additive noise, with implications beyond reconstruction tasks.
3 Data and Model Architecture
This section details the training and testing data, contamination models, and architectures used to study double descent.
@@ 3.1 Subspace Data Model
We revisited the model used in (Lupidi et al., 2023), which argued that “double descent does not occur in self-supervised settings.” We sampled $N$ i.i.d. Gaussian vectors of size $d$ (${\boldsymbol{z}}_{i}\sim\mathcal{N}(0,\mathbb{I}_{d})$), representing latent features and embedded them to a higher-dimensional space using ${\boldsymbol{H}}$ of size $D\times d$, ($D>d$, ${\boldsymbol{H}}_{ij}\sim\mathcal{N}(0,1)$). Our dataset differs from (Lupidi et al., 2023), and we explore four scenarios in our study:
@@
${\boldsymbol{x}}_{i}=\begin{cases}\beta{\boldsymbol{H}}{\boldsymbol{z}}_{i}+\boldsymbol{\epsilon}_{i},&\text{with probability }p,\\ \beta{\boldsymbol{H}}{\boldsymbol{z}}_{i},&\text{with probability }1-p,\end{cases}$ (1)
where $\boldsymbol{\epsilon}_{i}\sim\mathcal{N}(0,\mathbb{I}_{D})$ represents noise added to samples with probability $p$ and $\beta$ controls the SNR.
Feature Noise. We study the impact of noisy training features on the test loss curve by selecting the same set of $\lfloor D\cdot p\rfloor$ features to be noisy across all samples. This simulates $\lfloor D\cdot p\rfloor$ unreliable or noisy measuring tools.
A visualization of the data generation and SNR calculations for both sample and feature noise scenarios are presented in Appendix A, Figure 9 and Appendix B respectively.
Domain Shift. We analyze test loss behavior under domain shifts by projecting the test and train latent features using ${\boldsymbol{H}}$ and ${\boldsymbol{H}}^{{}^{\prime\prime}}$ respectively, with a shift modeled as ${\boldsymbol{H}}^{{}^{\prime\prime}}={\boldsymbol{H}}+s\cdot{\boldsymbol{H}}^{{}^{\prime}}$, where ${\boldsymbol{H}}_{ij}^{{}^{\prime}}\sim\mathcal{N}(0,1)$ adds perturbations, and $s>0$ controls the shift,
${\boldsymbol{x}}_{i}=\begin{cases}{\boldsymbol{H}}{\boldsymbol{z}}_{i},&\text{if }train,\\ {\boldsymbol{H}}^{{}^{\prime\prime}}{\boldsymbol{z}}_{i},&\text{if }test.\end{cases}$
@@
Anomalies. We examine how train-set anomalies affect the test loss curve. Normal samples are $\{\beta{\boldsymbol{H}}{\boldsymbol{z}}_{i}\}_{i=1}^{N}$, while anomalies follow $\mathcal{N}(0,\mathbb{I}_{D})$. The Signal-to-Anomaly Ratio (SAR), controlled by $\beta$, sets their magnitude ratio. We replace $p\cdot 100\%$ of normal samples with anomalies. This process is shown in Appendix A, Figure 11.
@@ 3.2 Single-Cell RNA Data
We used single-cell RNA data from (Tran et al., 2020) to demonstrate our findings in a challenging real-world setting. This dataset, which spans diverse domain shifts across laboratory environments and measurement technologies, is crucial for assessing domain shift impact on the test loss curve. Due to its real-world nature, we cannot control the shifts between the training (source) and test (target) datasets.
@@ 3.3 CelebA Data
To evaluate if the double descent phenomenon can impact crucial downstream tasks such as anomaly detection, we used the CelebA dataset (Han et al., 2022). Appendix A provides more details about our real-world datasets.
@@ 3.4 Fully-connected Neural Network (FCN)
We used FCNs (see Figure LABEL:fig:mlp_ae_architecture) to study double and triple descent across all mentioned data models. FCNs are widely used with tabular data and have recently demonstrated state-of-the-art results in various tasks (Gorishniy et al., 2024a, b; Shenkar & Wolf, 2022; Rauf et al., 2024; Svirsky & Lindenbaum, ). To demonstrate the generality of our result, we further demonstrate double descent using CNNs (see architecture in Appendix A, Figure LABEL:fig:cnn_ae_architecture) trained on image datasets are presented in Appendix F.3.
@@ 4 No Double Descent in Linear AEs
Notation and setup. Consider a general linear AE, $\Phi(\cdot;\boldsymbol{\theta}):\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$, which consists of $L\geq 2$ layers and is parameterized by $\boldsymbol{\theta}\triangleq$$\left\{{\boldsymbol{W}}_{1}^{d_{1}\times D},\ldots,{\boldsymbol{W}}_{L-1}^{d_{L-1}\times d_{L-2}},{\boldsymbol{W}}_{L}^{D\times d_{L-1}}\right\}$. Here, ${\boldsymbol{W}}_{i}$ represents the weights parameters associated with the $i$-th layer, for $i\in[L]$. For simplicity, we denote $\mathcal{D}=\{d_{i}\mid i\in[L]\}$ (including $d_{L}=d_{0}=D$). Accordingly, we let $m=\min_{i\in[L]}d_{i}$ to be the bottleneck of the autoencoder, namely, $\forall i\in[L],m\leq d_{i},$ and $m\leq D.$ Formally, the model $\Phi$ is given by, $\Phi({\boldsymbol{x}};\boldsymbol{\theta})={\boldsymbol{W}}_{L}{\boldsymbol{W}}_{L-1}\cdots{\boldsymbol{W}}_{2}{\boldsymbol{W}}_{1}{\boldsymbol{x}}$.
@@
$\frac{1}{N}\min_{\boldsymbol{\theta}}\sum_{i=1}^{N}\mathcal{L}(\boldsymbol{\theta};{\boldsymbol{x}}_{i})=\min_{\boldsymbol{\theta}}\frac{1}{N}\sum_{i=1}^{N}\left\|\Phi({\boldsymbol{x}}_{i};\boldsymbol{\theta})-{\boldsymbol{x}}_{i}\right\|_{2}^{2}.$
According to the formulation of noisy training samples in (1), let $\mathcal{P}_{{\boldsymbol{x}}}$ denote the distribution from which the ${\boldsymbol{x}}_{i}$’s are drawn, i.e., ${\boldsymbol{x}}_{i}\sim\mathcal{P}_{{\boldsymbol{x}}}$. Similarly, let $\mathcal{P}_{\tilde{{\boldsymbol{x}}}}$ represent the distribution from which the non-noisy evaluation samples $\tilde{{\boldsymbol{x}}}_{i}$’s are drawn (see the bottom part of (1)). Then, for the trained $\Phi({\boldsymbol{x}};\boldsymbol{\theta}^{opt}),$ where $\boldsymbol{\theta}^{opt}=\text{argmin}_{\boldsymbol{\theta}}\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(\boldsymbol{\theta};{\boldsymbol{x}}_{i})$, we formalize the generalization error (GE) on the distribution of the clean data as:
$\mathcal{G}(\mathcal{D};\boldsymbol{\theta}^{opt},\mathcal{P}_{\tilde{{\boldsymbol{x}}}})=\mathbb{E}_{\mathcal{P}_{\tilde{{\boldsymbol{x}}}}}\left[\|\Phi(\tilde{{\boldsymbol{x}}};\boldsymbol{\theta}^{opt})-\tilde{{\boldsymbol{x}}}\|_{2}^{2}\right].$
@@
Theoretical results. The following sentence shows that double descent does not occur in linear AEs. We use the above notations, where a subscript ($a$ or $b$) distinguishes between different AEs.
@@
Let ${\boldsymbol{X}}=({\boldsymbol{x}}_{1},{\boldsymbol{x}}_{2},\ldots,{\boldsymbol{x}}_{N})\in\mathbb{R}^{D\times N}$ be the training set, i.i.d. sampled as (1). Assume that the number of samples $N$ is large enough such that $\mathrm{rank}({\boldsymbol{X}})=D$ almost surely.
@@
Let $\Phi_{a}(\cdot;\boldsymbol{\theta}),\,\Phi_{b}(\cdot;\boldsymbol{\theta}):\mathbb{R}^{D}\to\mathbb{R}^{D}$ be two linear AEs, each with $L\geq 2$ layers and architectures defined by $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$, with bottleneck sizes $m_{a}$ and $m_{b}$, where $m_{a}<m_{b}\leq D$. Let $\boldsymbol{\theta}_{a}^{\text{opt}}$ and $\boldsymbol{\theta}_{b}^{\text{opt}}$ be the optimal parameters under assumption 4.1. Let $d$ denote the dimension of the support of $\mathcal{P}_{\tilde{{\boldsymbol{x}}}}$ (the dimension of the span of the samples $\tilde{{\boldsymbol{x}}}\sim\mathcal{P}_{\tilde{{\boldsymbol{x}}}}$). Then, the following holds,
$\mathcal{G}\left(\mathcal{D}_{b};\boldsymbol{\theta_{b}}^{\text{opt }},\mathcal{P}_{\tilde{{\boldsymbol{x}}}}\right)\leq\mathcal{G}\left(\mathcal{D}_{a};\boldsymbol{\theta_{a}}^{\text{opt }},\mathcal{P}_{\tilde{{\boldsymbol{x}}}}\right),$
@@
The proof of Theorem 4.2 can be found in Appendix C. This result implies that the GE (test loss) decreases as the bottleneck size $m$ increases. The theorem implies that the GE remains unchanged if the bottleneck dimension is preserved, even if the size of other layers increases.
@@
Figure 4: Subspace data model exhibits model-wise non-monotonic behavior for varying domain shifts. Appendix D, Figure 12 shows the train loss behavior.

### v3

@@ Abstract
The phenomenon of double descent has challenged the traditional bias-variance trade-off in supervised learning but remains unexplored in unsupervised learning, with some studies arguing for its absence. In this study, we first demonstrate analytically that double descent does not occur in linear unsupervised autoencoders (AEs). In contrast, we show for the first time that both double and triple descent can be observed with nonlinear AEs across various data models and architectural designs. We examine the effects of partial sample and feature noise and highlight the critical role of bottleneck size in shaping the double descent curve. Through extensive experiments on both synthetic and real datasets, we uncover model-wise, epoch-wise, and sample-wise double descent across several data types and architectures. Our findings indicate that over-parameterized models not only improve reconstruction but also enhance performance in downstream tasks such as anomaly detection and domain adaptation, highlighting their practical value in complex real-world scenarios.
@@ 1 Introduction
Figure 1: Demonstration of double descent phenomenon with unsupervised AEs for the “sample noise” scenario (Subsection 3.1). We present the test loss for varying epochs and hidden layer sizes.
In this study, we use AEs to explore double descent and its impact on key unsupervised tasks like domain adaptation, anomaly detection, and robustness to noisy data. AEs are widely used in unsupervised tasks such as denoising (Vincent et al., 2008; 2010), manifold learning (Wang et al., 2014; Duque et al., 2020), clustering (Song et al., 2013; Yang et al., 2019), anomaly detection (Sakurada & Yairi, 2014; Zhou & Paffenroth, 2017), feature selection (Han et al., 2018; Gong et al., 2022), domain adaptation (Deng et al., 2014; Yang et al., 2021), segmentation (Myronenko, 2019; Baur et al., 2021), and generative modeling (Kingma, 2013; Doersch, 2016), making them a prominent use case when studying double descent in unsupervised learning.
We start by analyzing linear AEs, whose lack of nonlinear activation functions allows for theoretical exploration. We prove and empirically confirm that double descent does not occur in linear unsupervised AEs. We then investigate nonlinear AEs and find that multiple descents emerge across different data models and architectures.
We provide extensive evidence showing the phenomenon in unsupervised nonlinear AEs trained on contaminated data, with “memorization” causing overfitting and performance degradation. However, we find that over-parameterized models can extract the true signal, leading to a second descent and improved performance.
@@
These findings provide critical insights into the dynamics of unsupervised learning and the interplay between model capacity, noise, and performance. The phenomenon is shown in Figure 1, illustrating the bias-variance tradeoff, the critical regime, and the second descent induced by over-parametrization of unsupervised AEs.
Our findings have important implications for key tasks in unsupervised learning. Specifically, we show that over-parameterized models adapt better to target domains when trained on source domains despite domain shifts. We also reveal non-monotonic performance in anomaly detection as model complexity increases, underscoring the practical relevance of our results, particularly when addressing outliers and domain shifts.
@@
(a) To the best of our knowledge, this is the first demonstration of model and epoch, and sample-wise double descent in a fully unsupervised setting. We analyze and contrast the behavior of linear and nonlinear AEs through analytical and empirical evaluations, highlighting their distinct dynamics.
@@ 2 Related Work
In this work, we study AEs’ unsupervised objective of minimizing reconstruction error, $\|\Phi({\bm{x}})-{\bm{x}}\|^{2}_{2}$, where $\bm{x}$} is the input and $\Phi({\bm{x}})$ is the AE’s output. This learning framework, which has not previously shown double descent, falls outside the settings explored in prior studies. We find that linear AEs lack double descent, while nonlinear AEs show double and triple descent, underscoring the role of nonlinearity. Our analysis spans model, epoch, and sample levels, exploring the impact of bottleneck size, hidden layers, noise magnitude, and sample contamination. We also demonstrate double descent in real-world unsupervised scenarios like domain shifts, anomalies, and additive noise, with implications beyond reconstruction tasks.
3 Data Model
This section details the training data, testing data, and contamination models used to study double descent.
@@ 3.1 Subspace Data Model
We revisited the model used in (Lupidi et al., 2023), which argued that “double descent does not occur in self-supervised settings.” We sampled $N$ i.i.d. Gaussian vectors of size $d$ (${\bm{z}}_{i}\sim\mathcal{N}(0,\mathbb{I}_{d})$), representing latent features and embedded them to a higher-dimensional space using $\bm{H}$} of size $D\times d$, ($D>d$, ${\bm{H}}_{ij}\sim\mathcal{N}(0,1)$). Our dataset differs from (Lupidi et al., 2023), and we explore four scenarios in our study:
@@
${\bm{x}}_{i}^{{D\times 1}}=\begin{cases}\beta{\bm{H}}{\bm{z}}_{i}+\bm{\epsilon}_{i},&\text{with probability }p,\\ \beta{\bm{H}}{\bm{z}}_{i},&\text{with probability }1-p,\end{cases}$ (1)
where $\bm{\epsilon}_{i}\sim\mathcal{N}(0,\mathbb{I}_{D})$ represents noise added to samples with probability $p$ and $\beta$ controls the SNR.
Feature Noise. We study the impact of noisy training features on the test loss curve by selecting the same $\lfloor D\cdot p\rfloor$ features to be noisy across all samples. This simulates $\lfloor D\cdot p\rfloor$ unreliable or noisy measuring tools.
When noise is added to the data, we quantify the ratio between signal and noise power using the SNR. A high SNR indicates high-quality (clean) data, whereas a low SNR corresponds to lower-quality (noisy) data, where the noise dominates. We report SNR values in decibels, defined as $\text{SNR [dB]}=20\cdot\log_{10}(\frac{E[\|\beta{\bm{H}}{\bm{z}}\|_{2}]}{E[\|\bm{\epsilon}\|_{2}]})$. To ensure the training data is sufficiently noisy, we consider cases where $\text{SNR}\leq 0$, meaning the noise power exceeds that of the signal. The SNR calculations for both sample and feature noise scenarios are presented in Appendix B, and a visualization of the data generation is illustrated in Appendix A, Figure 7.
Domain Shift. We analyze test loss behavior under domain shifts by projecting the test and train latent features using $\bm{H}$} and $\bm{H}$$^{{}^{\prime\prime}}$ respectively, with a shift modeled as ${\bm{H}}^{{}^{\prime\prime}}={\bm{H}}+s\cdot{\bm{H}}^{{}^{\prime}}$, where ${\bm{H}}_{ij}^{{}^{\prime}}\sim\mathcal{N}(0,1)$ adds perturbations, and $s>0$ controls the shift,
${\bm{x}}_{i}=\begin{cases}{\bm{H}}{\bm{z}}_{i},&\text{if }train,\\ {\bm{H}}^{{}^{\prime\prime}}{\bm{z}}_{i},&\text{if }test.\end{cases}$
@@
Anomalies. We examine how train-set anomalies affect the test loss curve. Normal samples are $\{\beta{\bm{H}}{\bm{z}}_{i}\}_{i=1}^{N}$, while anomalies follow $\mathcal{N}(0,\mathbb{I}_{D})$. The Signal-to-Anomaly Ratio (SAR), controlled by $\beta$, sets their magnitude ratio. We replace $p\cdot 100\%$ of normal samples with anomalies. Illustration is shown in Appendix A, Figure 9.
We additionally provide results for a nonlinear subspace data model across all the aforementioned contamination scenarios (i.e., sample and feature noise, domain shift, and anomalies) in Appendix F.4.
In the following subsections (i.e., 3.2, 3.3, 3.4), we present real-world, high-dimensional datasets from diverse domains, trained on different architectures. This setup allows us to demonstrate the generalization of the double descent phenomenon across both datasets and model types. Due to their high dimensionality, these datasets are well-suited for dimensionality reduction via AEs, as also explored in prior work (Eraslan et al., 2019), (Yan et al., 2016), (Wang et al., 2016), (Meng et al., 2017).
@@ 3.2 Single-Cell RNA Data
We used single-cell RNA data from (Tran et al., 2020) to demonstrate our findings in a challenging, high-dimensional real-world setting. The dataset was chosen due to its inherent domain shifts, which include 5 distinct domains (biological batches) stemming from differences in laboratory conditions and measurement technologies. This makes it particularly suitable for testing our claims about double descent under real-world distribution shifts, which are common in biological and medical data (Haghverdi et al., 2018; Shaham et al., 2017). Each batch represents 15 different cell types, where each cell (sample) in this dataset contains over 15,000 genes (features), making it a high-dimensional dataset. Due to the real-world nature, we cannot control the shifts between the training (source) and test (target) datasets. This dataset is used for sample noise, feature noise, and domain shift experiments.
@@ 3.3 CelebA Data
The CelebA dataset (Liu et al., 2015) was selected to evaluate the impact of model complexity on unsupervised anomaly detection performance (Han et al., 2022). CelebA provides a rich, real-world setting with data attributes, making it particularly suitable for testing whether model complexity translates into non-monotonic trends in downstream tasks such as anomaly detection.
3.4 MNIST and CIFAR-10
We employed the MNIST (LeCun et al., 1998) and CIFAR-10 (Krizhevsky et al., 2009) datasets to evaluate our findings on standard image benchmarks, demonstrating double descent under sample noise, feature noise, and domain shift scenarios. These datasets also support reproducibility across different model architectures and data domains.
For more information about the nonlinear data model, single-cell RNA, celebA, MNIST, and CIFAR-10 datasets, refer to Appendix A.
4 Model Architecture
This section details the model architectures used to study double descent.
@@ 4.1 Fully-connected Neural Network (FCN)
We used FCNs (see architecture in Figure LABEL:fig:mlp_ae_architecture) to study double and triple descent across all mentioned data models. FCNs are widely used with tabular data and have recently demonstrated state-of-the-art results in various tasks (Gorishniy et al., 2024a; b; Shenkar & Wolf, 2022; Rauf et al., 2024; Svirsky & Lindenbaum, ).
4.2 Convolutional Neural Network (CNN)
To illustrate the generality of our findings, we additionally present double descent phenomena using CNNs (architecture provided in Appendix A, Figure LABEL:fig:cnn_ae_architecture) trained on image datasets. A subset of the results is included in Subsection 6.1, while the majority, along with further experimental details, can be found in Appendix F.3. The results in the main paper for this architecture also include the training loss, whereas all other figures in the main paper, based on FCNs, show only test losses, with corresponding training losses provided in the Appendix. Across all experiments involving CNNs, the training loss consistently shows a monotonic decrease, similar to the other settings with FCNs.
@@ 5 No Double Descent in Linear AEs
Notation and setup. Consider a general linear AE, $\Phi(\cdot;\bm{\theta}):\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$, which consists of $L\geq 2$ layers and is parameterized by $\bm{\theta}\triangleq$$\left\{{\bm{W}}_{1}^{d_{1}\times D},\ldots,{\bm{W}}_{L-1}^{d_{L-1}\times d_{L-2}},{\bm{W}}_{L}^{D\times d_{L-1}}\right\}$. Here, ${\bm{W}}_{i}$ represents the weights parameters associated with the $i$-th layer, for $i\in[L]$. For simplicity, we denote $\mathcal{D}=\{d_{i}\mid i\in[L]\}$ (including $d_{L}=d_{0}=D$). Accordingly, we let $m=\min_{i\in[L]}d_{i}$ to be the bottleneck of the AE, namely, $\forall i\in[L],m\leq d_{i},$ and $m\leq D.$ Formally, the model $\Phi$ is given by, $\Phi({\bm{x}};\bm{\theta})={\bm{W}}_{L}{\bm{W}}_{L-1}\cdots{\bm{W}}_{2}{\bm{W}}_{1}{\bm{x}}$.
@@
$\frac{1}{N}\min_{\bm{\theta}}\sum_{i=1}^{N}\mathcal{L}(\bm{\theta};{\bm{x}}_{i})=\min_{\bm{\theta}}\frac{1}{N}\sum_{i=1}^{N}\left\|\Phi({\bm{x}}_{i};\bm{\theta})-{\bm{x}}_{i}\right\|_{2}^{2}.$
According to the formulation of noisy training samples in (1), let $\mathcal{P}_{{\bm{x}}}$ denote the distribution from which the ${\bm{x}}_{i}$’s are drawn, i.e., ${\bm{x}}_{i}\sim\mathcal{P}_{{\bm{x}}}$. Similarly, let $\mathcal{P}_{\tilde{{\bm{x}}}}$ represent the distribution from which the non-noisy evaluation samples $\tilde{{\bm{x}}}_{i}$’s are drawn (see the bottom part of (1)). Then, for the trained $\Phi({\bm{x}};\bm{\theta}^{opt}),$ where $\bm{\theta}^{opt}=\text{argmin}_{\bm{\theta}}\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(\bm{\theta};{\bm{x}}_{i})$, we formalize the generalization error ($\mathcal{G}$) on the distribution of the clean data as:
$\mathcal{G}(\mathcal{D};\bm{\theta}^{opt},\mathcal{P}_{\tilde{{\bm{x}}}})=\mathbb{E}_{\mathcal{P}_{\tilde{{\bm{x}}}}}\left[\|\Phi(\tilde{{\bm{x}}};\bm{\theta}^{opt})-\tilde{{\bm{x}}}\|_{2}^{2}\right].$
@@
Theoretical results. The following theorem shows that double descent does not occur in linear AEs. We use the above notations, where a subscript ($a$ or $b$) distinguishes between different AEs.
@@
Let ${\bm{X}}=({\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{N})\in\mathbb{R}^{D\times N}$ be the training set, i.i.d. sampled as (1). Assume that the number of samples $N$ is large enough such that $\mathrm{rank}({\bm{X}})=D$ almost surely.
@@
Let $\Phi_{a}(\cdot;\bm{\theta}),\,\Phi_{b}(\cdot;\bm{\theta}):\mathbb{R}^{D}\to\mathbb{R}^{D}$ be two linear AEs, each with $L\geq 2$ layers and architectures defined by $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$, with bottleneck sizes $m_{a}$ and $m_{b}$, where $m_{a}<m_{b}\leq D$. Let $\bm{\theta}_{a}^{\text{opt}}$ and $\bm{\theta}_{b}^{\text{opt}}$ be the optimal parameters under assumption 5.1. Let $d$ denote the dimension of the support of $\mathcal{P}_{\tilde{{\bm{x}}}}$ (the dimension of the span of the samples $\tilde{{\bm{x}}}\sim\mathcal{P}_{\tilde{{\bm{x}}}}$). Then, the following holds,
[…]
@@
Recall that the models are trained using noisy data sampled from $\mathcal{P}_{{\bm{x}}}$ and tested on clean data drawn from $\mathcal{P}_{\tilde{{\bm{x}}}}$ to see if there is noise memorization (high test loss) or signal learning (low test loss).
The proof of Theorem 5.2 can be found in Appendix C. This result shows that the $\mathcal{G}$ (test loss) decreases as the bottleneck size $m$ increases. Interestingly, it implies that the $\mathcal{G}$ remains unchanged if the bottleneck dimension is preserved, even if the size of other layers increases. Since the proof does not depend on the similarity between $\mathcal{P}_{{\bm{x}}}$ and $\mathcal{P}_{\tilde{{\bm{x}}}}$, the theorem holds for scenarios where $\mathcal{P}_{{\bm{x}}}\neq\mathcal{P}_{\tilde{{\bm{x}}}}$, including the cases of domain shifts and anomalies discussed in Subsection 3.1.